OpenStudy (anonymous):

if only 12.5% Mo remains after 200 hours, show that its half life is 66.7 hrs

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

ok its me so what is the question again cuz I handed in the wkst but I have the problem done

OpenStudy (anonymous):

the mummy one?

OpenStudy (anonymous):

your choice

OpenStudy (anonymous):

both? haha im really bad at chem

OpenStudy (anonymous):

ok we'll start with the one you just posted

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we are going label Ao= 1 because its 100% A= 12.5% or .125

OpenStudy (anonymous):

we have to solve for k first and we also know t=200

OpenStudy (anonymous):

now use the equation \[A =A_oe^(-kt)\]

OpenStudy (anonymous):

are ewe using A= Aoe -kt equation

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just plug in the numbers and write equation not doing any math

OpenStudy (anonymous):

now it take the ln.125=lne-k200 lne goes away so its just ln.125=-k200

OpenStudy (anonymous):

there isn't a negative sorry

OpenStudy (anonymous):

on the k200?

OpenStudy (anonymous):

with the last thing I said

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok...got that

OpenStudy (anonymous):

i have.... ln.125=k200

OpenStudy (anonymous):

what do you have now

OpenStudy (anonymous):

now divideln.125 by 200

OpenStudy (anonymous):

-.01039

OpenStudy (anonymous):

did you divide with a neg

OpenStudy (anonymous):

no....i did ln(.125) then i divided it by 200

OpenStudy (anonymous):

then you were suppose to keep the negative haha

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

then you have to use the half life equation ln2/k which is the value you just found to get the half life

OpenStudy (anonymous):

woah...i got it!

OpenStudy (anonymous):

yay!!!!!

OpenStudy (anonymous):

now if i can remember it for the test hah

OpenStudy (anonymous):

except you have to find 200hrs

OpenStudy (anonymous):

and you just work backwards

OpenStudy (anonymous):

are you ready for the mummy question

OpenStudy (anonymous):

what...no i dont it says show that he half like is 66.7 hrs and thats what i just did

OpenStudy (anonymous):

ohhh on the test you mean?

OpenStudy (anonymous):

should i just write it here or post it as a new question?

OpenStudy (anonymous):

yeah that's for the test and just write the question here

OpenStudy (anonymous):

a mummy is found to be 3940 years old by C dating. Using a half life of 5730 years for C, demonstrate that 62.1% of the C at the death of the person remains in his mummy

OpenStudy (anonymous):

ok so use the half life equation with ln2/t(1/2) =k

OpenStudy (anonymous):

what is the equation before math

OpenStudy (anonymous):

\[Ao \div 2= Aoe -kt1/2\]

OpenStudy (anonymous):

no use ln2/t t=the half life time

OpenStudy (anonymous):

the ln2=kt1/2

OpenStudy (anonymous):

no ln2/t=k

OpenStudy (anonymous):

you need to find k before anything

OpenStudy (anonymous):

isnt it ln2/t1/2=k

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I didn't want to confuse by adding the 1/2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so now what is t1/2

OpenStudy (anonymous):

5730 years

OpenStudy (anonymous):

yeah so solve for k now

OpenStudy (anonymous):

1.209x10^-4

OpenStudy (anonymous):

good now use the A=Aoe^-kt Ao=100% or 1and we know k and t is the other time now

OpenStudy (anonymous):

and were solving for A?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

use the e to the powet button

OpenStudy (anonymous):

then whats the 62.1%

OpenStudy (anonymous):

oh is that A, hahah

OpenStudy (anonymous):

A=e^-.4766 -----> A=.621 or 62.1%

OpenStudy (anonymous):

yea haha

OpenStudy (anonymous):

i got 50.01%

OpenStudy (anonymous):

did you get e^-.4766

OpenStudy (anonymous):

no mine was e^-.6927

OpenStudy (anonymous):

you multiplied (1.21x10^-4)x3940

OpenStudy (anonymous):

oh nno...i used 5730...oops

OpenStudy (anonymous):

yeah you have to go back to the other time

OpenStudy (anonymous):

ok lemme recheck

OpenStudy (anonymous):

ok got it...now is that how its worded on the test

OpenStudy (anonymous):

ok I just did it and I got it

OpenStudy (anonymous):

yeah that one is the same

OpenStudy (anonymous):

ok..and are the nuclear equations the same or does he change those

OpenStudy (anonymous):

there are some of the same and som enew ones

OpenStudy (anonymous):

I have to go do you have any last questions

OpenStudy (anonymous):

ok and the quarterly is all the same justt in multiple choice form?

OpenStudy (anonymous):

hyeah exactly the same

OpenStudy (anonymous):

ok thank you so much!!